SPECTROSCOPY OF LUMINESCENT CRYSTALS CONTAINING RARE EARTH ELEMENTS

MENG-LING CHEN, KWANG-HWA LII, and \underline{BOR-CHEN CHANG}, Department of Chemistry, National Central University, 300 Jhongda Road, Jhongli 32001, Taiwan.

We have studied the spectroscopy of luminescent crystals containing rare earth elements such as KEuGe_{2}O_{6},^{a} Cs_{3}EuSi_{6}O_{15},^{b} K_{4}[(UO_{2})Eu_{2}(Ge_{2}O_{7})_{2}], and R_{2}(C_{8}H_{10}O_{4})_{3} (R= Y, Tb, or Eu). The emission and excitation spectra of these compounds were recorded at ambient temperature. These spectra are consistent with the structures which were determined by single crystal X-ray diffraction. Crystals containing hybrid luminescent centers were also synthesized and interesting energy transfer mechanisms were observed. For example, dramatic luminescence quenching was found in KEu_{x}Nd_{1−x}Ge_{2}O_{6} (x= 0.98, 0.96, 0.94, and 0.84) as well as in Cs_{3}Eu_{0.98}Nd_{0.02}Si_{6}O_{15}, while different compositions of Y_{x}Eu_{y}Tb_{2−x−y}(C_{8}H_{10}O_{4})_{3} exhibit different emission colors. Emission lifetimes were also measured for these compounds, and the results shed light on the energy transfer mechanisms. Detailed results of our research will be presented.

aP.-L. Chen, P.-Y. Chiang, H.-C. Yeh, B.-C. Chang, and K.-H. Lii, Dalton Trans., 1721 (2008).

bM.-Y. Hung, Y.-H. Chen, B.-C. Chang, and K.-H. Lii, Chem. Mater. 17, 5743 (2005).